The length of the sides of a square was increased by 50%. by how many percent did the area of the square increased?
Answer by checkley77(12844)

LET X BE THE ORIGINAL SIDE.
LET 1.5X BE THE INCREASES SIDE.
X*X=X^2 FOR THE ORIGINAL SIZE.
1.5X*1.5X=2.25X^2
2.25X^2/X^2=2.25 IS THE INCREASE.